# Spatial Context-based Self-Supervised Learning for Handwritten Text Recognition

### v2

@@ 1 Introduction
This work was motivated by the relevance of spatial information in HTR tasks [25] together with the fact that, to our best knowledge, it has not been previously studied, and focuses on the adaptation of a reference set of such approaches to the HTR field. More precisely, we propose two alternatives that enable the use of these SSL strategies in HTR scenarios (see Fig. 1): (i) adapting the HTR task to the original SSL method, namely input adaptation, and (ii) proposing novel spatial context-based SSL methods devised to the characteristics of the HTR task, namely method adaptation. These proposals are comprehensively evaluated by considering several reference HTR corpora and are compared to the reference strategies followed in the field. The results obtained show that the proposed SSL strategies provide state-of-the-art recognition rates compared to reference SSL approaches with less computational requirements, while maintaining conceptual simplicity.
@@ 2.3 Spatial context-based methods
Despite being largely explored in classification frameworks, spatial context-based SSL is rarely exploited in image-based text recognition tasks. Nevertheless, the competitive results that are obtained by employing these strategies for classification tasks, together with the strong relationship between spatial information and HTR—owing to stroke and font style—suggest that, if adequately adapted, text recognition models should benefit from pre-training methods of this nature. Note that, our work not only seeks to understand and unravel the existing spatial context SSL methods, but also presents innovative methods that specifically leverage and capitalize on this spatial context, thus providing new perspectives and advancements within HTR.
@@ 6.2 Computational architectural requirements
Table 4: Table showing the input size, GFLOPs, and number of parameters for each method.
@@ 6.3 Strengths and limitations
Concerning the limitations, it has also been proven that dual-branch architectures or generative methods show a higher abstraction capacity in the representations by encapsulating more semantic information. In this regard, since the presented proposals exclusively focus on the visual information extracted from the stroke and font style, its applicability to other off-the-shelf tasks—e.g., text segmentation, writer retrieval or image enhancement—remains uncertain. Future works could explore the generalizability of the different SSL methods.

### v3

@@ 1 Introduction
This work was motivated by the relevance of spatial information in HTR tasks [11, 26] together with the fact that, to our best knowledge, it has not been previously studied, and focuses on the adaptation of a reference set of such approaches to the HTR field. More precisely, we propose two alternatives that enable the use of these SSL strategies in HTR scenarios (see Fig. 1): (i) adapting the HTR task to the original SSL method, namely input adaptation, and (ii) proposing novel spatial context-based SSL methods devised to the characteristics of the HTR task, namely method adaptation. These proposals are comprehensively evaluated by considering several reference HTR corpora and are compared to the reference strategies followed in the field. The results obtained show that the proposed SSL strategies provide state-of-the-art recognition rates compared to reference SSL approaches with less computational requirements, while maintaining conceptual simplicity.
@@ 2.3 Spatial context-based methods
Despite being largely explored in classification frameworks, spatial context-based SSL is rarely exploited in image-based text recognition tasks. Nevertheless, the competitive results that are obtained by employing these strategies for classification, together with the strong relationship between spatial information and HTR—owing to stroke and font style [11]—suggest that, if adequately adapted, text recognition models should benefit from pre-training methods of this nature. Then, while previous efforts on SSL for HTR have focused on contrastive or image generation frameworks, this work explores the spatial context-based paradigm. Note that, our work not only seeks to understand and unravel the existing spatial context SSL methods, but also presents innovative methods that specifically leverage and capitalize on this spatial context, thus providing new perspectives and advancements within HTR.
@@ 6.2 Computational architectural requirements
Table 4: Table showing the input size, GFLOPs, and number of parameters for each method. Symbol * denotes that the number of columns is dynamically adjusted to preserve the image aspect ratio.
@@ 6.3 Strengths and limitations
Concerning the limitations, it has also been proven that dual-branch architectures or generative methods show a higher abstraction capacity in the representations by encapsulating more semantic information. In this regard, since the presented proposals exclusively focus on the visual information extracted from the stroke and font style, its applicability to other off-the-shelf tasks—e.g., text segmentation, writer retrieval, image enhancement, multilingual or historical manuscripts—remains uncertain. Future works could explore the generalizability of the different SSL methods.